If the temperatur on the fahrenheit scale if F degrees and on Celcius scale is C degress, then C=5/9(F-32). What is the set of values of F if C is between 10 and 20?

Answer by alicealc(293)

C = 5/9 * (F - 32)

If C = 10, then
10 = 5/9 * (F- 32)
10 * 9/5 = F - 32
90/5 = F - 32
18 = F - 32
18 + 32 = F
50 = F

If C = 20, then
20 = 5/9 * (F- 32)
20 * 9/5 = F - 32
180/5 = F - 32
36 = F - 32
36 + 32 = F
68 = F

So, if C is between 10 and 20 then F is between 50 and 68